OpenStudy (anonymous):

solve for x x/4-4/x=0

OpenStudy (anonymous):

can you please rewrite the question

OpenStudy (anonymous):

x\[\frac{ x }{ 4 } - \frac{ 4 }{ x }\]

OpenStudy (anonymous):

= 0

OpenStudy (anonymous):

do you know how to cross multiply?

OpenStudy (anonymous):

yes...then i need to isolate x.

OpenStudy (anonymous):

ok you start with \[\frac{x}{4}-\frac{4}{x}=0\] move 4/x over so you can cross multiply \[\frac{x}{4}=\frac{4}{x}\] what do you get after you cross multiply???

OpenStudy (anonymous):



OpenStudy (anonymous):

ok, i was confused because of the = sign

OpenStudy (anonymous):

can you finish this on your own?

OpenStudy (anonymous):

yes thanks so much x=4 after all the math

OpenStudy (anonymous):

x=-4 also

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\sqrt{16}= \pm4\]

OpenStudy (anonymous):

why is that?